OpenStudy (anonymous):

solve sqrt 9x+55=x+5

jhonyy9 (jhonyy9):

squared both sides 9x+55=x2 +10x +25 x2 +x -30=0 (x+6)(x-5) so x1= -6 and x2 = 5